OpenStudy (anonymous):

Find the nth term for the following sequences (with index starting at n=1) SEE BELOW!

OpenStudy (anonymous):

(a) 8, -9/4, 10/9 (b)6/5, 10/9, 14/13

OpenStudy (anonymous):

I have that the difference between the denominators for A is 1, 3, 5. So i'm assuming the denominator follows the odd order, but thats all i have

OpenStudy (anonymous):

I know I generally say something is odd by (2K+1), but that gives me a first term of 3, if k starts at 1

OpenStudy (anonymous):

so far for A i have \[(-1)^{n+1}\frac{ n+7 }{ ? }\]

OpenStudy (anonymous):

the first part compensates for the alternating sign, and the numerator seems to fit. I am just not sure how to represent the denominators

OpenStudy (austinl):

Try this for part a, \(\displaystyle(-1)^{n+1}(\frac{n+7}{n^2})\)

OpenStudy (anonymous):

omg wow, duh! thanks

OpenStudy (austinl):

I believe that should get you the denominators that you are looking for with this sequence. Do you have any thoughts for part b?

OpenStudy (anonymous):

well for the denominators there is a difference of 4 for each one

OpenStudy (anonymous):

and for the denominator as well

OpenStudy (anonymous):

but i'm stuck from there

OpenStudy (austinl):

I am drawing a bit of a blank.... I can almost get one to work, but.... I am a bit rusty with these :P

OpenStudy (anonymous):

No problem. I think I will head to my teacher's office tomorrow. Thanks a bunch though!

OpenStudy (austinl):

No problem, @amistre might be able to help. He is more well rounded than me.

OpenStudy (austinl):

@amistre64 forgot the 64.....

OpenStudy (anonymous):

thanks i'll do that